OpenStudy (anonymous):

Which of the following ordered pairs represents the solution to the system given below? x − 4y = 7 5x + 9y = 6 A) 3,-1 B) 3,1 C) 1,-3 D) -1, -3

OpenStudy (anonymous):

@Mehek14

OpenStudy (ashes.boom):

you can put the coordinates into the equation to check them or you can solve for one variable and then use that to get the other

OpenStudy (ashes.boom):

add 4y to the first equation on both sides to get x by itself then you have x = 4y + 7

Mehek (mehek14):

if you graph it, the solution would be where they intersect

OpenStudy (anonymous):

Thanks

OpenStudy (ashes.boom):

then put that into the equation for x 5(4y + 7)+9y = 6 distribute 20y + 35 + 9y = 6 combine like terms 29y + 35 = 6 - 35 -35 29y = -29 divide by 29 to get y alone /29 /29 y = -1

OpenStudy (ashes.boom):

then in x = 4y + 7 put in -1 for y x = 4(-1) + 7 x = -4 + 7 x = 3

OpenStudy (ashes.boom):

the solution is (3, -1)

OpenStudy (anonymous):

thanks @Ashes.boom

OpenStudy (ashes.boom):

yea your welcome :)